Two forces equal to p and 2p act on particle when the first force is increased by 120N and the second force is doubled the direction of the resultant remains the same find the value of p?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
for first two forces,
Resultant1 = root(p^2 + 4p^2 + 4p^2 *costhetha).......................(1)
after increasing forces,
Resultant2 = root( (p+120)^2  + (4p)^2 + 2*4p*(p+120)costhetha).................(2)
and the direcn. is same therefore,
tanthetha = p/Resultant1...............(3)
and
 
tanthetha = (p+120)/Resualtant2.................(4)
solve the four eqns.
and get the values of R1 R2 thetha and p.